# The NOvA Collaboration

### v1

@@ Abstract
The flux of cosmic ray muons at the Earth’s surface exhibits seasonal variations due to changes in the temperature of the atmosphere affecting the production and decay of mesons in the upper atmosphere. Using data collected by the NOvA Near Detector during 2018–2022, we studied the seasonal pattern in the multiple-muon event rate. The data confirm an anticorrelation between the multiple-muon event rate and effective atmospheric temperature, consistent across all the years of data. Previous analyses from MINOS and NOvA saw a similar anticorrelation but did not include an explanation. We find that this anticorrelation is driven by altitude–geometry effects as the average muon production height changes with the season. This has been checked with a CORSIKA cosmic ray simulation package by varying atmospheric parameters, and provides an explanation to a longstanding discrepancy between the seasonal phases of single and multiple-muon events.
@@ I Introduction
The flux of muons produced in cosmic ray air showers and detected underground exhibits a well-known seasonal variation. This effect arises due to the competition between the decay and interaction of secondary mesons ($\pi,K$) produced in the primary cosmic-ray interaction. In the upper atmosphere, where most muons originate, higher temperatures in summer cause the atmosphere to expand, leading to lower density and a greater probability for mesons to decay into muons before interacting. Conversely, in winter, the atmosphere contracts, resulting in higher density, which increases the likelihood of mesons interacting before they can decay. As a result, the muon flux reaches a maximum in summer and a minimum in winter. In the past, this seasonal effect has been observed in multiple experiments Barrett et al. (1952); Sherman (1954); Castagnoli and Dodero (1967); Fenton et al. (1961); Y. Andreyev et al. () (Baksan Collaboration); Ambrosio et al. (1997); Bouchta ; Bellini et al. (2012); M. Selvi et al. () (LVD Collaboration); Desiati et al. ; S. Osprey et al. (2009) (MINOS Collaboration); Adamson et al. (2010, 2014); T. Abrahao et al. (2017) (Double Chooz Collaboration); F. P. An et al. (2018) (Daya Bay Collaboration). For our case, since the NOvA near detector is underground, the majority of the cosmic muons produced in the atmosphere do not reach the detector as the energy of the primary cosmic rays falls exponentially ($\sim E^{-2.7}$). Only a muon with an energy $E_{\mu}>\sec{\theta_{zen}}\times$ 50 GeV can reach the detector, where $\theta_{zen}$ is the muon zenith angle.
Large neutrino detectors are capable of detecting multiple-muon events, where several nearly parallel muon tracks from a single cosmic ray air shower arrive at the detector. While the seasonal effect of single muons is well understood, the behavior of multiple-muon events presented an intriguing anomaly. Previous studies, such as those from MINOS Adamson et al. (2015) and NOvA Acero et al. (2019), reported a winter maximum in observed multiple-muon rates, in direct contrast to the summer peak seen for single muons. To investigate this discrepancy, we conducted a detailed analysis of multiple-muon events using the NOvA Near Detector (ND). We concentrate on the size of the muon component of air showers compared to the size of our ND. Our approach incorporates a cosmic muon simulation using COsmic Ray SImulations for KAscade (CORSIKA) Heck et al. , along with four years of NOvA ND data. This study provides new insights into the longstanding puzzle of multiple-muon seasonal variations and offers an explanation through the altitude–geometry effect, described in Section III. Further sections describe the more recent NOvA ND data, temperature data, the CORSIKA simulation, cosine fits to the data, and a discussion of other seasonal effects.
II Previous data from MINOS and NOvA
@@
The MINOS experiment first reported an unexpected winter maximum in multiple-muon events underground, contrasting with the well-established summer peak observed for single muons. This effect was observed in both the MINOS ND Adamson et al. (2015), located at 225 meters water equivalent (mwe), and the MINOS Far Detector (FD) Adamson et al. (2015), located at 2100 mwe underground. In the MINOS FD, the seasonal behavior was found to depend on the observed spatial separation of muons in an event, with closely spaced muons exhibiting a winter maximum while widely separated muons followed the expected summer maximum behavior. The MINOS study proposed that secondary hadronic interactions in the atmosphere played an important role in this result, though a definitive explanation was not presented.
The NOvA ND Acero et al. (2019) ($E_{\mu}~>$ 50 GeV), located at the same depth as the MINOS ND but with a different detector design, confirmed the winter peak in multiple-muon events in a dataset spanning 2015–2017. The NOvA FD Acero et al. (2021) ($E_{\mu}~>$ 1.5 GeV), situated at the surface with minimal overburden, also observed a strong seasonal dependence in multiple-muon showers, with a 30% stronger winter maximum at higher multiplicities compared to the lowest multiplicities and a 20% stronger effect for showers arriving from near the horizon compared to the most vertical showers. The previous NOvA ND study Acero et al. (2019) found that the strength of the seasonal variation, which was measured in three different ways (Rayleigh Power, correlation coefficient, and cosine fit), did not clearly depend on zenith angle, muon separation, or muon angular separation, but did noticeably depend on muon multiplicity.
The MINOS study explored four possible explanations for this effect Adamson et al. (2015), including the altitude–geometry effect, but ultimately discarded it based on the difference between winter and summer in the separation among muons in the detector.
@@
In this paper, we extend the analysis of Ref. Acero et al. (2019) with four years of NOvA ND data, explain the seasonal variation of multiple-muon events with the altitude–geometry effect, and check that explanation with a CORSIKA Heck et al. based simulation of cosmic rays in the NOvA ND.
III The Altitude–Geometry Effect
The altitude–geometry explanation of the multiple-muon seasonal effect is based on the fact that a multiple-muon shower that originates higher in the atmosphere spreads out further than one originating lower. Thus, more muons from the shower miss a finite-size underground detector in summer than in winter. Although there would be more multiple-muon events in the summer, fewer of them would be detected as multiple-muon events because they originate higher in the atmosphere. This is the main explanation for the observed seasonal pattern of multiple-muon events with the MINOS Adamson et al. (2015) and NOvA Acero et al. (2019, 2021) detectors.
The total mass of the atmosphere remains approximately constant throughout the year, but seasonal temperature differences cause vertical expansion in summer and contraction in winter. Under the assumption of an isothermal atmosphere, the ideal gas law $PV=nRT$ implies that a $\pm$2% change in absolute temperature results in a corresponding $\pm$2% shift in altitude for a given pressure level. For a shower origin at fixed pressure, the increased altitude in summer leads to a broader lateral spread of muons at the detector. In contrast, during winter, the lower altitude results in a more concentrated footprint as shown in Figure 1 Kokoulin et al. (2016). As a result, a pair of muons that would be identified as a multiple-muon event in winter may be detected as a single-muon event in summer, simply due to their larger separation at the detector level.
This is precisely what we observed in the CORSIKA simulation in section VII. Moreover, calculations by members of the IceCube collaboration based on the altitude–geometry effect seemed to explain the MINOS and NOvA multiple-muon seasonal data Gaisser and Verpoest (2021a, b). In Figure 2, it can be observed that the birth altitude of muons from multiple muon events is higher in summer than in winter. The mean altitude increases in the summer compared to the winter. As a result, the separation among muons in multiple-muon events also increases in summer compared to winter as seen in Figure 3.
In contrast, the single muon rate remains unaffected. For every single muon that misses a finite-sized detector because it starts at a higher altitude, there is another single muon that now hits the detector. There is no detector geometry effect for single muons: a consequence of muons starting at a higher altitude in the summer affects the observed muon multiplicity in a finite detector, but does not change the muon flux.
Figure 2: The production altitude distribution of multiple-muon events from CORSIKA with muon threshold energy of 50 GeV. Three winter months (Dec–Feb) and three summer months (Jun–Aug) are compared. Vertical lines represent the mean muon birth altitude for summer (red dashed line) and winter (black solid line).
Figure 3: Muon separation distribution of multiple-muon events from CORSIKA with muon threshold energy of 50 GeV. Three winter months (Dec–Feb) and three summer months (Jun–Aug) are compared. Vertical lines represent the mean muon separation for summer (red dashed line) and winter (black solid line).
It is instructive to consider the typical transverse momentum of particles in hadronic showers, and how that impacts the separation distribution of muons reaching a detector. The opening angle between a primary pion and the original cosmic ray varies as $\sim(p_{T}/p_{l})$, where $p_{T}$ is the transverse momentum of the pion and $p_{l}\sim E_{\pi}$ is the longitudinal momentum, relative to the direction of the primary cosmic ray. A typical separation for muons from two pions in the primary interaction would be $\sqrt{2}\times A\times(p_{T}/E_{min})$, where $A$ is the altitude where the pions decay and $E_{min}$ is larger for deeper detectors due to muon energy loss in the overburden. For a typical hadronic transverse momentum of $300\text{\,}\mathrm{M}\mathrm{e}\mathrm{V}$ for two $60\text{\,}\mathrm{G}\mathrm{e}\mathrm{V}$ pions decaying at an altitude of $20\text{\,}\mathrm{k}\mathrm{m}$, one would expect a typical muon separation of $140\text{\,}\mathrm{m}$ at the NOvA ND. This is much larger than the size of the NOvA ND or MINOS ND or MINOS FD, so the altitude–geometry effect will decrease the number of observed multiple-muon events in the summer. In contrast, if the typical muon separation is smaller than the detector, which can be the case for deep detectors with TeV muon thresholds, the altitude–geometry effect is minimal. For example, the surface area of the MACRO experiment Ambrosio et al. (2002) is $76\text{\,}\mathrm{m}$$\times$$12\text{\,}\mathrm{m}$, but the threshold energy for an atmospheric muon to reach the detector is $1.3\text{\,}\mathrm{T}\mathrm{e}\mathrm{V}$. At this energy, the average separation among muons is approximately $8\text{\,}\mathrm{m}$, which is smaller than the size of the MACRO detector. Hence, MACRO observed a summer maximum for multiple-muon events Ronga (2017), just as for single muon events. However, in the MINOS FD, the typical muon separation is comparable to the size of the detector. This helps to explain why the observed muon separation in the detector determines whether there is a summer maximum or a winter maximum.
Every detector has a unique size, but there is a straightforward relationship between the depth of an underground detector and the minimum muon energy required to reach that detector. A detector at or near the surface can detect muons of a few GeV, and the size of an extended air shower can exceed a kilometer. At the MINOS FD, the muon energy threshold is near $0.7\text{\,}\mathrm{T}\mathrm{e}\mathrm{V}$, and the typical muon separation is a few meters. In other words, as detector depth increases, the muons that survive are coming from a smaller and smaller area of the core of the shower.
The role of summer/winter differences in the origin altitude of muons in an air shower, and its effect on shower containment in a finite detector, has been considered before by MINOS Adamson et al. (2015). However, the MINOS paper’s conclusion, discarding the effect, was based on the measured muon separation distributions in the near and far detectors without taking into account that the measured distribution does not reflect the true separation if the size of the muon shower is larger than the detector. The altitude–geometry effect was considered responsible for the seasonal results of the DECOR experiment Yurina et al. (2019), which measured muons close to the surface of the earth. For those muons, with an energy typically of a few GeV, other effects such as pressure changes are important, as well as muon decay, which is discussed further in Section VIII.
IV COSMIC MUON DATA IN THE NOvA NEAR DETECTOR
@@
In each cell, the light emitted by the scintillator is collected and transported to an avalanche photodiode (APD) through a wavelength shifting fiber. The APD converts the light to an electronic signal, digitized by the front-end electronics, and recorded as a hit if the signal exceeds a preset threshold value.
@@
Cosmic muons in the NOvA ND are collected using a data-driven activity trigger. This trigger initiates data collection when there are at least 10 hits in detector planes, requiring hits on at least 3 planes in each of the two views and a minimum of 8 planes in total. Additionally, hits must be recorded in at least 5 planes out of 6 consecutive planes.
@@
Collected hits are reconstructed using the Hough Transform Ballard (1981) method which is widely used for line reconstruction. All possible reconstructed tracks in two dimensions (2D) are stored. These 2D tracks from each plane are merged to create a 3D track.
@@
In order to ensure a clean sample of cosmic ray induced multiple muon events, three selection cuts are applied to the tracks. First, the track’s start and endpoint should be less than 50 cm from the detector surface. This fiducial cut eliminates a number of events in which an electron emerged from the rock above the cavern, nearly parallel to an accompanying muon, and made a short track. This cut also eliminates contained and partially contained muon events. Each muon track is required to cross at least 10 planes. Some single muon events that occur closely spaced in time are recorded as multiple-muon events. Hence, the time between the first and any other tracks of a multiple-muon event is limited to a maximum of $100\text{\,}\mathrm{n}\mathrm{s}$. This eliminates those falsely recorded as multiple-muon events. The fiducial cut removes 16.9% of the total number of events considered in the analysis. The selection criteria on the number of planes further eliminate 16.2% of the total events, and finally the timing cut removes an additional 0.04% of the events.
@@
The track multiplicity distribution of selected events from the ND data is shown in Figure 4. The maximum multiplicity observed in the ND data is 10. The distribution of cosine of the zenith angle, representing the angle between the muon track and the vertical upward direction, is illustrated in Figure 5. The distribution exhibits good agreement between data and Monte Carlo simulation.
@@
Figure 5: The distribution of the cosine of the angle between the track and the vertical upward direction (zenith angle). The ratio of Data over the CORSIKA MC prediction is shown in the lower panel with statistical uncertainty.
The multiple-muon event rate in the data is obtained by fitting the distribution of time difference ($\Delta t$) between pairs of consecutive events. It eliminates the effects of any data acquisition glitches, such as brief pauses resulting in data loss, or bursts of noise. The time differences for such cases appear in the outliers of the distribution while the core of the distribution is unaffected. The distribution is fitted with the Erlang distribution function,
$R[\lambda,x\equiv\log_{10}(\Delta t)]=Ae^{-\lambda 10^{x}}10^{x}\lambda.$ (1)
Here, $\Delta t$ is the time difference between consecutive events, $\lambda$ is the event rate, and $A$ is a normalization constant. The fitted distribution is shown in Figure 6 in the range 1.5 to 4.7, and the $\chi^{2}$/NDF from the fits as a function of time are shown in Figure 7. This is an improvement over the previous method used in the NOvA ND analysis Acero et al. (2019) that calculated the rate by dividing the total number of muons by the total exposure time.

### v2

@@ Abstract
The flux of cosmic ray muons at the Earth’s surface exhibits seasonal variations due to changes in the temperature of the atmosphere affecting the production and decay of mesons in the upper atmosphere. Using 1473 live days of data collected by the NuMI Off-axis $\nu_{e}$ Appearance (NOvA) Near Detector during 2018–2022, we studied the seasonal pattern in the multiple-muon event rate. The data confirm an anticorrelation between the multiple-muon event rate and effective atmospheric temperature, consistent across all the years of data. Previous analyses from MINOS and NOvA saw a similar anticorrelation but did not include an explanation. We find that this anticorrelation is driven by altitude–geometry effects as the average muon production height changes with the season. This has been studied with a CORSIKA cosmic ray simulation package by varying atmospheric parameters, and provides an explanation to a longstanding discrepancy between the seasonal phases of single and multiple-muon events.
@@ I Introduction
The flux of muons produced in cosmic ray air showers and detected underground exhibits a well-known seasonal variation. This effect arises due to the competition between the decay and interaction of secondary mesons ($\pi,K$) produced in the primary cosmic-ray interaction. In the upper atmosphere, where most muons originate, higher temperatures in summer cause the atmosphere to expand, leading to lower density and a greater probability for mesons to decay into muons before interacting. Conversely, in winter, the atmosphere contracts, resulting in higher density, which increases the likelihood of mesons interacting before they can decay. As a result, the muon flux reaches a maximum in summer and a minimum in winter. In the past, this seasonal effect has been observed in multiple experiments Barrett et al. (1952); Sherman (1954); Castagnoli and Dodero (1967); Fenton et al. (1961); Y. Andreyev et al. () (Baksan Collaboration); Ambrosio et al. (1997); Bouchta ; Bellini et al. (2012); M. Selvi et al. () (LVD Collaboration); Desiati et al. ; S. Osprey et al. (2009) (MINOS Collaboration); Adamson et al. (2010, 2014); T. Abrahao et al. (2017) (Double Chooz Collaboration); F. P. An et al. (2018) (Daya Bay Collaboration). For our case, since the NuMI Off-axis $\nu_{e}$ Appearance (NOvA) Near Detector (ND) is underground, the majority of the cosmic muons produced in the atmosphere do not reach the detector as the energy of the primary cosmic rays falls exponentially ($\sim E^{-2.7}$). Only a muon with an energy $E_{\mu}>\sec{\theta_{zen}}\times$ 54 GeV can reach the detector, where $\theta_{zen}$ is the muon zenith angle.
@@
Large neutrino detectors are capable of detecting multiple-muon events, where several nearly parallel muon tracks from a single cosmic ray air shower arrive at the detector. While the seasonal effect of single muons is well understood, the behavior of multiple-muon events presented an intriguing anomaly. The MINOS experiment first reported an unexpected winter maximum in multiple-muon events underground, contrasting with the well-established summer peak observed for single muons. This effect was observed in both the MINOS ND Adamson et al. (2015), located at 225 meters water equivalent (mwe), and the MINOS Far Detector (FD) Adamson et al. (2015), located at 2100 mwe underground. The MINOS study considered possible explanations, and favored the idea that secondary hadronic interactions in the atmosphere played an important role. The NOvA ND Acero et al. (2019) ($E_{\mu}~>$ 54 GeV), located at the same depth as the MINOS ND but with a different detector design, confirmed the winter peak in multiple-muon events in a dataset spanning 2015–2017. To investigate this discrepancy, we conducted a detailed analysis of multiple-muon events using the NOvA ND. We concentrate on the size of the muon component of air showers compared to the size of our ND.
@@
In this paper, we extend the analysis of Ref. Acero et al. (2019) using approximately four years of NOvA ND data (April 30, 2018 to May 11, 2022, corresponding to 1473 live days over four years). We explain the seasonal variation of multiple-muon events with the altitude-geometry effect, and check that explanation with a CORSIKA Heck et al. based simulation of cosmic rays in the NOvA ND.
In Sec. II, there are details of the NOvA ND. Section III describes the simulation and modeling by season of cosmic muons with CORSIKA. In Sec. IV we summarize how we obtained temperature data. Next, Sec. V presents the analysis of new multiple-muon data and an improved event rate calculation. Then, Sec. VI explains the observed seasonal behavior of the MINOS and NOvA multiple-muon rate as due to the altitude–geometry effect. A method to determine whether this effect switches the phase of underground detectors as a function of detector size and depth is given.
II The NOvA Near Detector
@@
The basic detector element is a polyvinyl chloride (PVC) cell of cross section $3.9\text{\,}\mathrm{cm}$ $\times$ 6.6 $6\text{\,}\mathrm{cm}$ and length $3.9\text{\,}\mathrm{m}$, filled with liquid scintillator. A plane is formed by arranging these cells side by side. The planes are arranged alternately in horizontal and vertical orientations to enable reconstruction of a 3D track.
@@
Scintillation light produced in each cell is collected by a wavelength-shifting fiber and transmitted to an avalanche photodiode (APD). The APD converts the light into an electronic signal, which is digitized by front-end electronics. Signals above a preset threshold are recorded as detector hits.
III Cosmic muon simulation
Previous analyses of MINOS and NOvA did not compare their data with a cosmic simulation that included seasonal effects. For the first time, we used the CORSIKA simulation with seasonal effects incorporated in it. Cosmic muons are simulated with CORSIKA Heck et al. v7.7410, using GHEISHA for low-energy and QGSJET for high-energy hadronic interactions. Primary particles are generated between 200 and $10^{6}$ GeV covering the angular domain of $0^{0}$ to $90^{0}$ for zenith angles and $0^{0}$ to $360^{0}$ for azimuthal angles with a power-law energy spectrum of index 2.7-2.7 Navas et al. (2024). Each muon in a multiple-muon event is required to have an energy above $54\text{\,}\mathrm{GeV}$, the minimum needed to traverse the overburden above the NOvA ND. Muons incident at larger zenith angles require additional energy due to the longer rock depth. The surviving energy after propagation through rock of density $\rho$ and thickness $d$ is calculated using
$E_{\mu}=(E_{\mu,0}+\epsilon)e^{-bd\rho}-\epsilon,$ (1)
where $E_{\mu,0}$ is the initial energy, $b=4\times 10^{-6}~\text{cm}^{2}/\text{g}$, and $\epsilon=500$ GeV are standard-rock parameters Zyla et al. (2020). Only muons with $E_{\mu}>0$ are retained.
CORSIKA models atmospheric density with a five-layer parametrization. For the first four layers the mass overburden $X(h)$ decreases exponentially with height $h$ (from sea level),
$X(h)=a_{i}+b_{i}e^{-h/c_{i}},$ (2)
while in the fifth layer, the dependence is linear,
$X(h)=a_{5}-b_{5}h/c_{5}.$ (3)
Since there is no default profile for Fermilab, we extracted site-specific vertical atmospheric profiles from GDAS National Centers for Environmental Prediction (2024) (NCEP) for 2017 and fitted them to this parametrization to obtain the coefficients $a_{i}$, $b_{i}$, and $c_{i}$.
Proton showers are generated with these atmospheric parameters and the simulation settings above. Cosmic muons from each shower are distributed uniformly across an extended $xz$ (horizontal surface where $z$ axis is along the neutrino beam direction) plane encompassing the detector ($x\in[-32,32]$ m, $z\in[-30,46]$ m). However, the NOvA ND spans $x\in[-2,2]$ m and $z\in[0,16]$ m; the extended surface ensures inclusion of $99\%$ of muons intersecting the detector volume. For multiple-muon events, the first muon is placed randomly on this plane, while the others retain their relative spacings and orientations from CORSIKA. Showers producing at least one muon in the ND are propagated through the detector using GEANT4 Agostinelli et al. (2003); Allison et al. (2006, 2016). Upward-going neutrino-induced muons are not considered.
Seasonal effects are studied with two complementary approaches. First, we consider all muons in a CORSIKA event with sufficient energy to reach the ND depth, corresponding to an “infinite detector.” Second, we simulate detector response with the standard NOvA ND software Acero et al. (2020). The simulated multiple-muon rate for an infinite detector shows a summer maximum as displayed in Fig. 1, opposite that seen in multiple-muon measurements from data. These simulated infinite-detector events were passed through the finite ND using GEANT4 simulation, and the same analysis cuts described in Ref. Acero et al. (2019) and later in Sec. V were applied. As a result, the time of the seasonal maximum shifted from summer to winter, as shown in Fig. 2. This result from our simulation is consistent with previous observations (NOvA Acero et al. (2019) and MINOS Adamson et al. (2015)). The contrast between Fig. 1 and  2 strongly suggested that the difference was due to a geometrical effect related to the size of the detector, and not a result of differences in the meson content of the hadronic shower.
Figure 1: Seasonal variation of multiple-muon rate for an infinite detector at the depth of the NOvA ND from CORSIKA initiated by primary proton cosmic rays. With detector size not included, the summer peak from atmospheric density effects on the muon rate is seen. The solid curve shows a cosine fit.
Figure 2: Seasonal variation of multiple-muon rate for an $4\text{\,}\mathrm{m}$ by $15.9\text{\,}\mathrm{m}$ detector at the depth of the NOvA ND from CORSIKA initiated by the primary proton cosmic rays. The solid curve shows a cosine fit. With a detector size less than the average of the muon shower at this depth ($105\text{\,}\mathrm{m}$), the summer peak has turned into a summer deficit because individual muons in the shower miss the detector, causing the event to not be classified as a multiple muon event. Error bars represent statistical uncertainties.
The CORSIKA implementation does not fully capture the absolute magnitude of single-muon seasonal variation, which is well understood theoretically and experimentally Grashorn et al. (2010). Nevertheless, the atmospheric density profile provides a reasonable basis for the relative seasonal differences in hadronic shower development. For the present analysis, only the comparison of the relative phase and modulation strength between single- and multiple-muon events in a fixed detector is used.
IV Temperature data
Atmospheric temperature data are taken from ERA5 Hersbach et al. (2020), the fifth-generation atmospheric reanalysis produced by the European Centre for Medium-Range Weather Forecasts (ECMWF) Dee et al. (2011). They provide temperature readings at 37 pressure levels, distributed nonuniformly from 1 to 1000 hPa. Temperatures are extracted from ECMWF within the grid area ranging from (41N, 87W) to (42.5N, 88.5W), which corresponds to the NOvA ND location.
$T_{e\!f\!f}=\frac{\sum_{i=1}^{37}{T(P_{i})[W_{\pi}(P_{i})+W_{K}(P_{i})]\Delta P_{i}}}{\sum_{i=1}^{n}[W_{\pi}(P_{i})+W_{K}(P_{i})]\Delta P_{i}}.$ (4)
Here, $T(P_{i})$ is the temperature at pressure level $i$ ($P_{i}$), and $W_{K}(P_{i})$ and $W_{\pi}(P_{i})$ are the weights responsible for the contribution as a function of pressure (altitude) for those $K$ and $\pi$ decays that make muons. The effective temperature is calculated using cross sections and lifetime values from Grashorn et al. (2010), along with temperature values obtained at 37 pressure levels from ECMWF. Figure 3 shows the variation of ECMWF temperature data of the atmosphere over the Fermilab area, averaged over the four years 2018–2022, as a function of altitude and pressure. It also shows the total weight $W_{\pi}+W_{K}$ used in Eq. 4.
Figure 3: Variation of average temperature with altitude and total weight function used to calculate effective temperature. This weight corresponds to the fraction of cosmic ray primaries left at a given height: for example, about half have already interacted at a height of $18\text{\,}\mathrm{km}$.
V COSMIC MUON DATA IN THE NOvA NEAR DETECTOR
@@
Cosmic muons in the NOvA ND are collected using a data-driven activity trigger. This trigger initiates data collection when there are at least ten hits in detector planes, requiring hits on at least three planes in each of the two views ($xz$ and $yz$) and a minimum of eight planes in total. Additionally, hits must be recorded in at least five planes out of six consecutive planes.
@@
Collected hits are reconstructed using the Hough transform Ballard (1981) method which is widely used for line reconstruction. All possible reconstructed tracks in 2D are stored. These 2D tracks from each plane are merged to create a 3D track.
@@
To obtain a clean sample of cosmic-ray–induced multiple-muon events, additional selection criteria are imposed. First, a fiducial cut requires both the start and end points of each track to be within $50\text{\,}\mathrm{cm}$ of the detector surface. This removes events in which electrons, produced nearly parallel to muons in the surrounding rock, generate short tracks in the detector, as well as contained or partially contained muon events. Second, each muon track must cross at least 10 planes to suppress short, poorly reconstructed tracks. Finally, to reduce contamination from single muon events occurring in close temporal proximity, the time difference between the first and any additional tracks in a multiple-muon event is required to be less than $100\text{\,}\mathrm{n}\mathrm{s}$. The fiducial requirement rejects 16.9% of events, the plane requirement removes 16.2%, and the timing cut discards an additional 0.04%.
@@
The track multiplicity distribution of selected events from the ND data is shown in Fig. 4. The maximum multiplicity observed in the ND data is 10. The distribution of cosine of the zenith angle, representing the angle between the muon track and the vertical upward direction, is illustrated in Fig. 5. The distribution exhibits good agreement between data and Monte Carlo simulation.
@@
Figure 5: Distribution of the cosine of the angle between the track and the vertical upward direction (zenith angle). The ratio of data over the CORSIKA MC prediction is shown in the lower panel with statistical uncertainty.
There are two differences between the multiple muon data in this analysis and that in  Acero et al. (2019). First, four years of new data are presented here. Second, the first analysis obtained a rate by simply dividing the number of selected events in a time period by the elapsed time covered by a data file. Both terms have uncertainties that, although small, can be difficult to quantify. For example, events that are not cosmic ray induced multiple muons might increase the event count, or unaccounted for dead time in the data acquisition might decrease the live time. Furthermore, the NOvA data is not stored in time order on disk, resulting in events that “belonged” in one file being stored in another, changing counts in both files. This analysis uses a fit based on sorted arrival time differences to eliminate these issues. Hence, we first reordered all of the multiple-muons based on their arrival time. Then rates are obtained by performing a fit to the distribution of time difference ($\Delta t$) between consecutive multiple-muon events with an Erlang’s probability density function (PDF) Forbes et al. (2011). The analytical form of the Erlang PDF,
$f(\Delta t;k,\lambda)=\frac{\lambda^{k}{\Delta t}^{k-1}e^{-\lambda\Delta t}}{(k-1)!}$ (5)
where $k$ is called the shape parameter and $\lambda$ is the rate parameter. The shape parameter $k$ is equal to 1 for the two consecutive multiple-muon events. So, putting $k=1$ for consecutive events, the PDF becomes
$f_{t}(\Delta t)=\lambda e^{-\lambda\Delta t},\qquad\Delta t\geq 0.$ (6)
Now under the transformation $x=\log_{10}(\Delta t)$, the Eq. 6 takes the form
$f_{x}(x)=A\,\lambda\,10^{x}\,e^{-\lambda 10^{x}},$ (7)
where $A$ is a normalization factor. This expression is the functional form used to extract the multiple-muon event rate from the $\log_{10}(\Delta t)$ spectrum. Since $\Delta t$ is computed in $ms$, $\lambda$ is obtained in $ms^{-1}$, which is then converted into $Hz$. A representative fit for a random week is shown in Fig. 6, with the quality of fits across all weekly bins summarized in Fig. 7.